OpenStudy (anonymous):

in triangle ABC, the sum of the measures of angle A and angle B is 47 degrees. what is the measure of angle c

OpenStudy (ryan123345):

133

OpenStudy (ryan123345):

just subtract 47 minus 180

jimthompson5910 (jim_thompson5910):

all angles must add to 180 degrees so a+b+c = 180 (a+b)+c = 180 47+c = 180 solve for c

OpenStudy (ryan123345):

the triangle sum theorem says that very triangle add up to 180. so u just subtract c:

OpenStudy (anonymous):

thankyou!

OpenStudy (ryan123345):

your welcome :)